Shift Summary for day shift, 30/july/99

Ashkan_Alavi-Harati (alavi@fnal.gov)
Fri, 30 Jul 1999 15:29:08 -0500

Shift Summary for day shift, 30/july/99

On Shift: Bruce Winstein, Ashkan Alavi

Hours of beam delivered at all intensities: 1.0
Hours of beam delivered at nominal intensity: 1.0
Hours of E832 data (special runs not included): 1.0
Hours of special runs (laser scans, muon runs...): 0.75

Accounting of lost time: a vacuum failure in the accelerator
followed by comEd power reduction.

========================================================================
Runs ended this shift:
Run 14062 BEAM832 12.7M events, 18587 vac 00's
Run 14063 LASER_SCAN

Tapes brought to FCC this shift:
(including tapes for online split)
QKC195-197, QKS195-197, QKK195-197, QKB68.

========================================================================
Accesses during shift: none

Two: see below.

========================================================================
Overview of shift:
Smooth E832 datataking before the power reduction shut down.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14062 252 08:00 Beam is gone, no message on the screen

08:25 Called the control room, there is a vacuum problem
in the accelerator. No estimate.

08:38 MCR called, there will be no beam before shut down

285 08:40 Stop the run.

14063 09:00 Laser Scan starts.

09:30 MC and AA made an access to measure the z
positions of the regenerator and DC1-2 with the laser system, but the
monitor on the PC was bad. Doug was notified.

10:15 Laser scan finished.

10:30 JH, AA, BW and a guest of BW's made an access.
Bruce to give his friend a tour, and Jason to look at the ET system.
Jason verified the SDAQ temperature reading. That digital thermometer is
reading 69 while the others are above 75. However, it is not a problem.
He also cycled the power for ET to fix the threshold problem. Thresholds
on all boards would get stuck at 69GeV, causing an error on the serious
alarms.

10:30 Ed installed new CSQ6 constants which are valid
for runs 14065-99999 (laser reference didn't get
updated).

11:45 Bruce cleaned all the tape drives.

========================================================================
Follow up Notes on Previous Shift

SDAQ button on Etot-temp is still read. The read out is 0.69 V instead
of 0.80 (corresponding to 80 degrees). Jason H verified this downstairs
on the digital thermometers. It doesn't seem to be any worries for now.
The red button can be ignored.

========================================================================
Items to Follow up on Next Shift